# InversionView: A General-Purpose Method for Reading Information from Neural Activations

### v3

@@ 3.5 Factual Recall
To test whether InversionView can be applied to larger language models, we explore how GPT-2 XL (1.5B parameters) performs the task of recalling factual associations. In this case study, our intention is not to provide a full interpretation of the computations performed to solve this task, which we deem out of scope for this paper. Instead, we show that InversionView produces interpretable results on larger models by focusing on a relatively small set of important attention heads in upper layers. The decoder model in this case study is based on GPT-2 Medium, because we expect a more complex inverse mapping from activation to inputs to be learned. We observe the resulting $\epsilon$-preimage can express high-level knowledge (Figure 38-43), and sometimes can predict the failure of the model (Appendix H.6). Using InversionView, we again shed light on the underlying mechanism of the model. We present detailed findings in Appendix H.
4 Discussion
@@
Traditional probing [e.g. 2, 6] trains a supervised classifier to assess how much information about a variable of interest is encoded in an activation site. It requires a hypothesis in advance and is thus inherently limited to hypotheses conceived a priori by the researcher. InversionView, on the other hand, helps researchers form hypotheses without any need for prior guesses, and allows fine-grained per-activation interpretation. Inspecting attention patterns [e.g. 12] is a traditional approach to inferring information flow, and we have drawn on it in our analyses. More recently, path patching [50, 26] causally identifies paths along which information flows. In contrast to these approaches, InversionView aims at aiding generation of hypotheses about what information is contained in activations and passed along. While the information flow provides an upper bound on the information passed along by tracing back to the input token, it is insufficient for determining how information is processed and abstracted. For instance, in Section 3.2, occurrences of the target character are causally connected to $a^{0,0}_{tc}$, which then connects to $a^{1,0}_{:}$ (direct or mediated by MLP layer 0). Without looking at encoded information, we only know that the information in these paths is related to the occurrences of the target character, but not whether it is their identity, positions, count, etc. More generally, when a component reads a component that itself has read from multiple components, connectivity does not tell us which pieces of information are passed on. In addition, Direct Logit Attribution (DLA) [50, 23, 40] is also frequently used for inspecting the output information of transformer’s components. We argue that DLA is only suitable for studying model components that directly affect model’s final output. For those components whose effect is mediated by other components, their output information is meant to be read by a downstream component, thus not necessarily visible when projecting to the vocabulary space. We provide further discussion in Appendix I.
The samples produced by InversionView can be easily fed into LLMs for automated interpretation. We show a proof of concept by using Claude 3 to interpret the model trained for 3-digit addition. See results in Table LABEL:tab:auto-inter. Overall, we find that the interpretation given by the LLM reflects the main information in almost all cases of the addition task. Despite some flaws, the outcome is informative in general, suggesting this as a promising direction for further speeding up hypothesis generation.
@@
Limitations.
@@
5 Related Work
Trained probing classifiers are arguably the most common method for uncovering information from activations [e.g. 2, 5, 4, 51, 48, 32, 33]. More recent methods obtain insights by projecting representations into the vocabulary space [40, 7, 41, 30, 24], patching them into an LLM [25], or by decomposing them into interpretable sparse feature vectors [10, 47, 14]. Dar et al. [16] directly interprets model parameters by projecting them into the embedding space. Another important area of research is uncovering information flow. Geva et al. [24] intervene on attention weights to study information flow. Activation patching, a causal intervention method, can be used to study the causal effect of an activation on the output, and can help localize where information is stored [36, 45], or find alignment between a high-level causal model and inner states of a neural model [21, 22, 53]. While activation patching affects all downstream computation in the model, path patching [50, 26] restricts effects for downstream components in order to identify the path along which information flows [50, 13, 27, 34]. Recent work has started using LLMs to generate interpretations [8, 10].
@@ 6 Conclusion
We present InversionView, an effective method for decoding information from neural activations. By applying it in four case studies—character counting, IOI, 3-digit addition, and factual recall—we showcase how it can reveal various types of information, thus facilitating reverse-engineering of algorithm implemented by neural networks. Moreover, we compare it with other interpretability methods and show its unique advantages. We also show that the results given by InversionView can in principle be interpreted automatically by LLMs, which opens up possibilities for a more automated workflow. This paper only explores a fraction of the opportunities this method offers. Future work could apply it to subspaces of residual stream (which can probably produce more insights), to larger models, or to different modalities such as vision.
7 Acknowledgements
@@ A.2 Sampling with Decoder Model
In Section 2.2, we mentioned that the distribution $p(\mathbf{x}|\mathbf{z}^{q})$ is modeled by the decoder. Strictly speaking, $p(\mathbf{x}|\mathbf{z}^{q})$ represents the data distribution in the $\epsilon$-preimage defined by $\epsilon=0$. For example, when the probed model is using causal masking, and a certain activation is relevant to all previous context (by non-zero attention weights), then $p(\mathbf{x}|\mathbf{z}^{q})$ is the distribution over those inputs that share the same previous context (i.e., they have same prefix). This requires that the decoder can distinguish any tiny difference in activation and decode the full information (imagine a token attended with 0.0001 attention weight). Such a decoder must be very powerful and perhaps trained without any regularization. But in practice, the decoder is a continuous function of activation and tiny changes in activation are not perceivable by the decoder. We observe that the decoder rarely generates the sample that lies at the same point (producing the same activation) as the query input in vector space, instead it usually generates samples that are in the neighborhood of the query input. Because we need to observe the whole neighborhood of the query input and prevent samples from being too concentrated, we adjust the sampling temperature to control how concentrated they are. Importantly, even if the decoder is too powerful and can always recover the same activation, we can still obtain the neighborhood by adding random noise to the query activation before giving it to the decoder. This motivates decoding with temperature and noise, as described in the next paragraph.
@@ Appendix B Experimental Verification of Completeness
In Section 2.2, we described that an ideal strategy for obtaining samples in the $\epsilon$-preimage satisfies two desiderata: it only provides samples that are indeed within the $\epsilon$-preimage (Correctness), and it provides all such samples at reasonable probability (Completeness). As further described there, we can directly ensure Correctness by evaluating $D(\cdot)$ for every sample. Ensuring completeness is more challenging, due to the exponential size of the input space; the most general approach is to design counterexamples not satisfying a hypothesis about the content of the $\epsilon$-preimage, and verifying that $D(\cdot)$ is indeed large.
@@ D.2 More Examples of InversionView
See Figures 13 and 14. Here, we show results similar to Figure 2 for other query inputs.
@@
Figure 15: Results of activation patching for model trained on character counting task. Same figure as 3(a) with intermediate steps of calculation shown using line plot. Note that the gray lines correspond to the y-axis on the right. In contrast examples, only one character differs. $LD$ stands for logit difference between the original count and the count in the contrast example. $LD_{pch}$ and $LD_{orig}$ correspond to the $LD$ with and without patching, respectively. Top: We patch activations cumulatively from left to right, flipping the sign of $LD$. The “none” on the left end of x-axis denotes the starting point, i.e., nothing is patched. Bottom: We patch from right to left. Similarly, “none” on the right end of x-axis denotes the starting point.
@@ E.4 Choice of Distance Metric in IOI
In Figure 22 and 23 we show two examples of using normalized Euclidean distance as the distance metric. Readers can see more examples on our web application. We see that the samples generated by the decoder tend to have larger distances under normalized Euclidean distance. Nonetheless, we can still see the top-ranked samples show meaningful commonalities – indeed if we set $\epsilon=0.4$, we obtain the same interpretation as the one we obtain based on Figure 3(b) and 18.
@@ F.2 More Examples of InversionView
See Figures 25 to 28.555We have also tried value-weighted attention pattern [31]; it makes a negligible difference, so we always show the original attention pattern in our paper.
@@ F.5 Causal Intervention: Details and Full Results
Activation patching results for other cases are shown in Figures 31 to 35.
@@ H.6 More Examples of InversionView
Appendix I Notes on Attention, Path Patching and DLA

### v4

@@ 3.4 Factual Recall
To test whether InversionView can be applied to larger language models, we explore how GPT-2 XL (1.5B parameters) performs the task of recalling factual associations. In this case study, our intention is not to provide a full interpretation of the computations performed to solve this task, which we deem out of scope for this paper. Instead, we show that InversionView produces interpretable results on larger models by focusing on a relatively small set of important attention heads in upper layers. The decoder model in this case study is based on GPT-2 Medium, because we expect a more complex inverse mapping from activation to inputs to be learned. We observe the resulting $\epsilon$-preimage can express high-level knowledge (Figure 39-44), and sometimes can predict the failure of the model (Appendix H.6). Using InversionView, we again shed light on the underlying mechanism of the model. We present detailed findings in Appendix H.
4 Discussion and Related Work
Comparison with other Interpretation Methods
@@ 4 Discussion and Related Work
Supervised probing classifiers, assessing how much information about a variable of interest is encoded in an activation site, are arguably the most common method for uncovering information from activations [e.g. 2, 6, 5, 4, 55, 52, 33, 34]. It requires a hypothesis in advance and is thus inherently limited to hypotheses conceived a priori by the researcher. InversionView, on the other hand, helps researchers form hypotheses without any need for prior guesses, and allows fine-grained per-activation interpretation. Inspecting attention patterns [e.g. 12] is a traditional approach to inferring information flow, and we have drawn on it in our analyses. More recently, path patching [54, 26, 13, 28, 35] causally identifies paths along which information flows. While the information flow provides an upper bound on the information passed along by tracing back to the input token, it is insufficient for determining how information is processed and abstracted. For instance, in Section 3.1, occurrences of the target character are causally connected to $a^{0,0}_{tc}$, which then connects to $a^{1,0}_{:}$ (direct or mediated by MLP layer 0). Without looking at encoded information, we only know that the information in these paths is related to the occurrences of the target character, but not whether it is their identity, positions, count, etc. More generally, when a component reads a component that itself has read from multiple components, connectivity does not tell us which pieces of information are passed on. Similar considerations apply to other intervention methods. Geva et al. [24] intervene on attention weights to study information flow. Activation patching, a causal intervention method, can be used to study the causal effect of an activation on the output, and can help localize where information is stored [37, 49], or find alignment between a high-level causal model and inner states of a neural model [21, 22, 57]. Many recent works obtain insights about information content by projecting representations or parameters into the vocabulary space [42, 7, 44, 31, 54, 23, 24, 16]. This technique is sometimes referred to as Direct Logit Attribution (DLA). We argue that DLA is only suitable for studying model components that directly affect model’s final output. For those components whose effect is mediated by other components, their output information is meant to be read by a downstream component, thus not necessarily visible when projecting to the vocabulary space. We provide further discussion in Appendix I. Generalizing this approach, Ghandeharioun et al. [25] patch activations into an LLM. Another line of recent research [10, 51, 14] decomposes activations into interpretable features using sparse autoencoders.
Some other interpretation methods also generate in input space, but differ from InversionView in goals and methods. This includes feature visualization [43, 41], adversarial or counterfactual example generation [27, 59, 46, 45], and GAN inversion methods [58]. We discuss the similarities and differences of these works compared to InversionView in Appendix I.4.
@@
Transformer Circuits for Arithmetic
Automated Interpretation for InversionView
Recent work has started using LLMs to generate interpretations [8, 10]. The samples produced by InversionView can be easily fed into LLMs for automated interpretation. We show a proof of concept by using Claude 3 to interpret the model trained for 3-digit addition. See results in Table LABEL:tab:auto-inter. The LLM-provided interpretation reflects the main information in almost all cases of the addition task. Despite some flaws, the outcome is informative in general, suggesting this as a promising direction for further speeding up hypothesis generation.
Limitations
@@ 5 Conclusion
We present InversionView, an effective method for decoding information from neural activations. In four case studies—character counting, IOI, 3-digit addition, and factual recall—we showcase how it can reveal various types of information, thus facilitating reverse-engineering of algorithm implemented by neural networks. Moreover, we compare it with other interpretability methods and show its unique advantages. We also show that the results given by InversionView can in principle be interpreted automatically by LLMs, which opens up possibilities for a more automated workflow. This paper only explores a fraction of the opportunities this method offers. Future work could apply it to subspaces of residual stream, to larger models, or to different modalities such as vision.
Acknowledgements
@@ A.2 Sampling with Decoder Model
In Section 2, we mentioned that the distribution $p(\mathbf{x}|\mathbf{z}^{q})$ is modeled by the decoder. Strictly speaking, $p(\mathbf{x}|\mathbf{z}^{q})$ represents the data distribution in the $\epsilon$-preimage defined by $\epsilon=0$. For example, when the probed model is using causal masking, and a certain activation is relevant to all previous context (by non-zero attention weights), then $p(\mathbf{x}|\mathbf{z}^{q})$ is the distribution over those inputs that share the same previous context (i.e., they have same prefix). This requires that the decoder can distinguish any tiny difference in activation and decode the full information (imagine a token attended with 0.0001 attention weight). Such a decoder must be very powerful and perhaps trained without any regularization. But in practice, the decoder is a continuous function of activation and tiny changes in activation are not perceivable by the decoder. We observe that the decoder rarely generates the sample that lies at the same point (producing the same activation) as the query input in vector space, instead it usually generates samples that are in the neighborhood of the query input. Because we need to observe the whole neighborhood of the query input and prevent samples from being too concentrated, we adjust the sampling temperature to control how concentrated they are. Importantly, even if the decoder is too powerful and can always recover the same activation, we can still obtain the neighborhood by adding random noise to the query activation before giving it to the decoder. This motivates decoding with temperature and noise, as described in the next paragraph.
@@ Appendix B Experimental Verification of Completeness
In Section 2, we described that an ideal strategy for obtaining samples in the $\epsilon$-preimage satisfies two desiderata: it only provides samples that are indeed within the $\epsilon$-preimage (Correctness), and it provides all such samples at reasonable probability (Completeness). As further described there, we can directly ensure Correctness by evaluating $D(\cdot)$ for every sample. Ensuring completeness is more challenging, due to the exponential size of the input space; the most general approach is to design counterexamples not satisfying a hypothesis about the content of the $\epsilon$-preimage, and verifying that $D(\cdot)$ is indeed large.
@@ Appendix C Decoder Model
The overall training and sampling pipelines are shown in Figure 2. In this section, we describe the architecture of the decoder model in detail.
@@ D.2 More Examples of InversionView
See Figures 14 and 15. Here, we show results similar to Figure 3 for other query inputs.
@@
Figure 16: Results of activation patching for model trained on character counting task. Same figure as 4(a) with intermediate steps of calculation shown using line plot. Note that the gray lines correspond to the y-axis on the right. In contrast examples, only one character differs. $LD$ stands for logit difference between the original count and the count in the contrast example. $LD_{pch}$ and $LD_{orig}$ correspond to the $LD$ with and without patching, respectively. Top: We patch activations cumulatively from left to right, flipping the sign of $LD$. The “none” on the left end of x-axis denotes the starting point, i.e., nothing is patched. Bottom: We patch from right to left. Similarly, “none” on the right end of x-axis denotes the starting point.
@@ E.4 Choice of Distance Metric in IOI
In Figure 23 and 24 we show two examples of using normalized Euclidean distance as the distance metric. Readers can see more examples on our web application. We see that the samples generated by the decoder tend to have larger distances under normalized Euclidean distance. Nonetheless, we can still see the top-ranked samples show meaningful commonalities – indeed if we set $\epsilon=0.4$, we obtain the same interpretation as the one we obtain based on Figure 4(b) and 19.
@@ F.2 More Examples of InversionView
See Figures 26 to 29.555We have also tried value-weighted attention pattern [32]; it makes a negligible difference, so we always show the original attention pattern in our paper.
@@ F.5 Causal Intervention: Details and Full Results
Activation patching results for other cases are shown in Figures 32 to 36.
@@ H.6 More Examples of InversionView
Appendix I Notes on Attention, Path Patching, DLA and others
@@ I.3 Additional Discussion about DLA
I.4 Methods Generating in Input Space
Feature visualization [43, 41] generates inputs maximally activating a certain neural network unit, and interprets an individual neural network unit (e.g., neurons) to understand its general role across inputs, while InversionView interprets specific values of inner representations, by finding inputs that result in the same vector. When the input changes, the value and thus the interpretation may change. Adversarial or counterfactual example generation methods [27, 59, 46, 45] generate input that is similar to the original input but results in different outcome. Some of them are also used to explain the model. While similar in input space, the adversarial/counterfactual input is likely to be quite different in internal representation space, leading to a different output. In contrast, we are interested in how different inputs in input space are represented very similarly in internal representation space. Similar to InversionView, GAN inversion methods [58] also study the mapping between input space and representation space, with a focus on interpreting the semantics of GAN’s latent space and manipulating generation.